U.S. Robotics (USR) has a current (and target) capital structure of 80 percent common equity and 20 percent debt. The beta for USR is 1.1. USR is evaluating an investment in a totally new line of business. The new investment has an expected internal rate of return of 25 percent.
USR wishes to evaluate this investment proposal. If the investment is made, USR intends to finance the project with the same capital structure as its current business. USR’s marginal tax rate is 30 percent. USR has identified three firms that are primarily in the line of business into which USR proposes expanding. Their average beta is 1.4, and their average capital structure is 40 percent common equity and 60 percent debt. The marginal tax rate for these three firms averages 40 percent. The risk-free rate is 7 percent, and the expected market risk premium is 8.8 percent. Should USR undertake the project? Round your answer to one decimal place.

The project should be accepted for any after-tax cost of debt of  % or -Select-higherless

please use fin. calculator
Answer #1

First lets delverage the beta of the similar businesses

= Beta / [ 1 + ( 1 - tax rate ) * debt / equity ]

= 1.40 * / [ 1 + ( 1-0.40) * 0.6/0.40 ] = 0.7368

Now lets releverage with the capital structure for USR

Levered beta = unlevered beta *[ 1 + ( 1 - tax rate ) * debt / equity ]  

= 0.7368 * [ 1 + ( 1-0.3) * 0.2/0.8]

= 1

Required rate of return as per CAPM = risk free rate + ( market return - risk free rate )* beta

= 7 + 8.8 * 1 = 15.8%

The required rate of return is lower than the IRR of the project hence the project can be accepted.